1. Two charges, +3.00mC and - 2.00mC, are separated by a distance of 5.00cm. What is the magnitude of the force they exert on each other? Is this repulsive of attractive?



2. An alpha particle moves 0.003m along the direction of an electric field of magnitude 1.00x104 N/C. An alpha particle has a charge of +2e and a mass of 6.68x10-²7 kg and e is equal to 1.60x10-¹9


A. What is the work done on the alpha particle?


B. If it was initially at rest, what is it velocity after moving 0.003m

1
Expert's answer
2022-05-15T17:02:32-0400

1.

"F=k\frac{q_1q_2}{r^2}"

"F=9*10^9*\frac{3.0*10^{-3}*2.0*10^{-3}}{0.05^2}=2.2*10^7\:\rm N"

Attractive.

2.

"W=Fq=qEd\\
=2*1.6*10^{-19}*1.0*10^{4}*0.003=9.6*10^{-18}\rm J"